OpenStudy (anonymous):

What is the 25th term of the geometric sequence that starts with .01 and has a ratio of 2?

OpenStudy (anonymous):

tn=ar^(n-1) here n=25 a=0.01 r=2 find t25

OpenStudy (anonymous):

the answer will be 167772.16 that is (0.01)*[(2)^(25-1)] = 0.01*(2^24) = 0.01*16777216 = 167772.16